# Benchmarking Misuse Mitigation Against Covert Adversaries

### v1

@@ 3.2 Evaluating LLM attacks on BSD questions
We next measure the effectiveness of existing attacks across a range of target models. In this section, we measure the accuracy across fifty harmful BSD multi-choice questions.
Baselines. To obtain strong misuse uplift baselines, we evaluate several attacks [^2] Because each attack we consider was primarily designed for open-ended generation tasks, we tailor them to make them compatible with multi-choice questions. For the PAIR, Adversarial Reasoning, and Crescendo, the method is applied only to the question, while the answer choices remain unchanged and are appended after the attacking prompt. We also modify the Adaptive Attack to use a multiple-choice template.. First, we consider three single-turn jailbreaks—simple adaptive attacks [25], PAIR [2], and adversarial reasoning attacks [21]—which attempt to obtain an answer via a single model interaction. We also use Crescendo [26], a multi-turn jailbreak, and decomposition attacks [18, 19], as described in Section 2. Moreover, we include a new decomposition attack variant described in detail in Section 5. Each attack is run for five epochs (when applicable) with a budget of twelve calls to the strong model per task. Each row denotes a different strong model, and we use Qwen2.5-7B as (a) the attacker for PAIR and adversarial reasoning attacks and (b) the weak model for decomposition attacks.
Results. As shown in Table 1, Claude-3.5 Sonnet and o3 yield the lowest accuracies among all strong targets. This mirrors Figure 1, where the same models exhibit the highest direct-query refusal rates. For both of these models, decomposition attacks are the strongest performing attacks. The advantage of decomposition attacks arises because the individual sub-queries evade detection, whereas jailbreaking prompts might be blocked by guardrails. For the other targets, whose safety filters are less effective, jailbreaking prompts are more likely to go undetected. We further investigate the detectability of decomposition and jailbreaking prompts in Section 4.
Table 1: Misuse rate for BSD of attacks on various strong models. The performance of our decomposition pipeline on misuse uplift significantly increases when the decomposer is fine-tuned to produce better sub-queries but still lacks the knowledge to solve the malicious task. These numbers show that the o3-mini model is highly prone to misuse.
@@
Adaptive PAIR Adversarial Reasoning Crescendo Decomposition Attack (theirs) Decomposition Attack (ours)
Claude-3.5-Sonnet 27.3 35.3 46.7 20.7 41.6 46.0
Claude-3.7-Sonnet 67.3 62.7 65.3 52.7 52.8 65.6
GPT-4o 42.0 64.7 70.7 65.3 68.4 74.0
o3-mini 84.7 84.7 84.0 86.1 82.0 81.2
o3 32.7 46.0 56.7 53.3 52.0 68.8
o1-preview 56.0 74.0 76.0 54.0 60.4 64.0
@@ 4 Detectability and defense
Figure 4: Decompositions are harder to identify than jailbreaks per-input. (Left) Hidden-layer UMAP projection [46] of Llama-3-8B [47]. Jailbreaking prompts (from [2, 21]) fall near the refused BSD questions, whereas decompositions are mostly located near benign WMDP questions. (Middle/Right) Decompositions are significantly harder to classify than jailbreaks in two settings (see 4.1).
Real-world LLM misuse (see, e.g., [23, 48]) typically involves completing multiple tasks, many of which may appear benign in isolation. This threat model is in tension with standard defenses, which assume that a harmful request is confined to a single context window. Decomposition attacks [19, 18] exploit this oversight and systematically avoid detection by splitting a harmful task into benign subtasks. We find that decomposition attacks are much harder to detect than standard jailbreaks. In Figure 4 (left), we plot Llama-3-8B activations on refused prompts (questions from the BSD dataset) and answered questions (WMDP), along with jailbreaks and decomposition subtasks. This provides evidence that jailbreaks fall closer to the refused BSD questions, while decompositions designed to appear benign are grouped with the answered WMDP questions. In the following, we find that this holds more generally across different defenses.
Figure 5: Benign prompts push the precision of an adversarially trained Llama-Guard classifier to zero by ‘drowning out’ the queries from decomposition attacks (for BSD biology). Lighter colors denote a higher ratio of benign prompts to harmful prompts. (Top Row) Left: Classifying between PAIR [2] prompts on BSD questions vs PAIR on answered/benign WMDP questions. Right: classifying at the user level significantly improves results (Bottom row) For our online stateful defense, we find that including an in-context buffer of the most suspicious user prompts significantly increases the detectability of harmful users for jailbreaks (Left) and decompositions (right).
@@ 4.1 Detecting harmful inputs
We find that decompositions of harmful questions are considerably harder to detect than standard jailbreak prompts. Here, we frame detection as deciding whether a given decomposition or jailbreak prompt was generated based on a query from WMDP or from BSD. Because frontier models almost never refuse WMDP [39] questions (see Section 3), but consistently refuse BSD questions, we label WMDP-derived prompts as safe and BSD-derived prompts as unsafe.
@@
As a baseline, we use Llama-Guard-3-8B to classify each prompt as safe or unsafe. In Figure 4 (middle), we plot the ROC curve and observe that for decomposition prompts, performance is only marginally better than random (AUC $=0.60$). In contrast, the same detector is significantly more effective on jailbreak prompts (AUC $=0.71$).
Adversarial training against decomposition attacks.
We allow the defender knowledge of both the attacker model and the misuse domain by giving them 20% of both the safe and unsafe decomposition or jailbreak prompts. We then finetune Llama-Guard-3-8B with the new data and calibrate the model to the new class distribution with Platt scaling. Adversarial training raises the AUC of Llama-Guard-3-8B for the jailbreak prompts (AUC $=0.89$) and the decompositions (AUC $=0.76$).
An adaptive attack against classification.
Anticipating that the defender is tuning their classification threshold or adversarially training (see above), the attacker floods the defender with a large number of benign tasks. As the benign-to-malicious ratio increases (lighter traces in the left column of Figure 5), precision collapses: distinguishing the now-rarer harmful prompts without excessive refusal becomes extremely challenging.
@@ 4.2 Detecting harmful users
Input-output detectors struggle to detect decomposition attacks because they lack the context necessary to recognize a harmful task distributed into multiple benign queries (see Section 4.1). We therefore lift the task to the user level: the defender succeeds if they flag an attacker before they are able to complete a harmful task. This assumes the model provider can trace all queries back to a single user, even if they occur in different sessions.
@@
For every user we maintain the $m$ prompts with the highest prompt-wise harmful scores in a buffer. For a new query, the buffer (ordered from most to least harmful) is fed in-context to either a fine-tuned Llama-Guard-3-8B or Llama-3-70B, which returns harmful or benign. The user is flagged if harmful is returned and subsequent requests are denied. Note that the defender has more information in this setting: we assume the attacker is in the middle of a misuse attempt and that there are (potentially) decomposition subtasks or jailbreak attempts in the user state.
Evaluation protocol.
We randomly sample a benign or harmful user. We sample $k\in\{10,\ldots,400\}$ previous user queries. If the user is benign, then the label of all of these prompts is benign. If the user is unsafe, there is some probability (the benign-to-malicious ratio) that a prompt is harmful. We rebuild the buffer with per-prompt harmfulness scores at each draw, place the buffer in-context to the classifier, and prompt it to reason over the previous prompts in the buffer to identify patterns of misuse. We show precision-recall results for classifying over ‘users’ in Figure 5 (right).
Results. Detection scores improve in the stateful setting. Stronger models with better in-context learning perform better in this setting. While user-level buffering only improves AUC from 0.76 to 0.79 for the fine-tuned Llama-Guard-3-8B, for Llama-3-70B the AUC improves from 0.71 to 0.86 This trend is clear in the right column of Figure 5: buffering shifts the precision–recall curve upward.
Take-aways. Maintaining a small rolling ’evidence window’ of suspicious queries in-context for a language model catches attacks that prompt-level filters miss. Because the method only requires storing a few recent queries and an additional inference call, it is relatively inexpensive. However, as can be seen in Figure 5, the adaptive attack (An adaptive attack against classification.) can still drive the false-positive rate up significantly for all the classifiers we consider.
@@ 5 Scaling decomposition attacks
The success of a decomposition attack depends on the quality of generated sub-queries, which, in turn, depends on factors including the coarseness of the decomposition and how comprehensively they span the original task. We show that two approaches can improve the performance of decomposition attacks: increasing the number of sub-tasks and distilling the model performing the decomposition.
Decomposition coarseness. One approach to measuring the performance of decomposition attacks is to increase the number of sub-tasks. In Figure 6 (left), we use Mixtral-8x22B as the weak model and GPT-4.1 as the strong model. We find that accuracy consistently improves as the number of decompositions increases. We also include a weak-model-only baseline, which uses the weak model to generate the decomposition and to answer the decomposed questions. The results for this baseline indicate that performance does not improve as the number of sub-tasks increases, which suggests that the strong model’s responses drive performance improvement.
In contrast, on tasks for which the target is willing to answer directly, decomposition attacks harm performance. As Figure 6 (right) shows, on WMDP, the accuracy of the decomposition attack is lower than the strong model’s direct query accuracy. This suggests that in general, decomposition attacks are not an effective test-time scaling technique, and therefore that the gains noted in Figure 6 (left) are due to uplift from the strong model’s answers.
Distilling an unsafe decomposer model with only benign data. We improve decomposition quality through benign-only distillation. We fine-tune only the model performing the decomposition. We collect 700 teacher-generated decompositions for benign MMLU-auxiliary [42] prompts using o3-mini, then fine-tuned Qwen on these prompts. For the distillation data, we randomly choose the number of decompositions for a question to have 3, 6, or 12 parts. Restricting the process to benign prompts allows us to use stronger models for distillation regardless of their safety guardrails. As the last two columns of Table 1 show, except for o3-mini that routinely answers the direct queries, the results unanimously improve on other strong models with the fine-tuned model. Our method achieves the state-of-the-art on Claude-3.5, OpenAI-o3 and GPT-4o.
@@
Figure 6: (Left, BSD) In the refusal setting with BSD, decomposition accuracy steadily improves as the number of decompositions increases. The baseline gets no answer from the strong model (GPT4.1). (Right, WMDP) On a similar dataset where strong models do not refuse (WMDP-Bio), decomposition consistently underperforms direct querying, suggesting the success of decomposition scaling is not from general test-time compute scaling.
6 Broader impact
We build a dataset of biology and cybersecurity questions that violate the content restrictions of various frontier model providers. This is because realistic misuse uplift evaluations require questions that are both difficult and consistently refused. For example, we found that tasks that were not dual-use, e.g., difficult math questions [49] or standard jailbreaking behaviors [4] failed to yield any questions that met our difficult and refused criteria (Section 3). Given concerns around disseminating harmful information, we do not do a full release of the BSD generated questions.
This research studies and develops strong attacks to misuse language models. While presenting these methods could enable attackers, suppressing them would likely hurt progress on effective countermeasures. Consistent with security-through-transparency norms, we discuss both attacks and mitigation strategies (Sections 33.2). We maintain that the security benefits of empowering the research community outweigh the incremental risk of adversary adoption.
@@ 7 Conclusion
We introduce a evaluation framework for measuring misuse uplift and detectability. Whereas previous evaluations measure if an attack can elicit harm from a given model, our framework measures the extent to which a strong model aides in misuse. We construct a threat model with realistic affordances for both the attacker (the ability to use weaker models) and the defender (tracking user queries across independent user conversations to detect misuse across contexts). We find that decomposition attacks [18, 19] are a particularly effective attack in this setting, outperforming state-of-the-art single- and multi-turn jailbreaks. We develop a defense that mitigates misuse with stateful detectors that reason over many independent user inputs to detect clusters of harmful inputs, however we find that decomposition attacks can subvert such detectors.
Acknowledgments and Disclosure of Funding
We thank David Rein, Weiqiu You, Adam Karvonen, and Anton Xue for feedback on the paper. The work of DB, MS, GP and HH is supported by the NSF CAREER award CIF-1943064 and the DARPA SAFRON grant, under award number HR0011-25-3-0135. The work of EW, DB, and LS is supported by ARPA-H grant D24AC00253-00, NSF award CCF 2313010, and by a gift from AWS AI to ASSET (Penn Engineering Center on Trustworthy AI). DB acknowledges support from AISS on a recommendation by Open Philanthropy.

### v2

@@ 4.1 Evaluating LLM attacks on BSD questions
We next measure the effectiveness of existing attacks across a range of target models. We measure the accuracy across fifty harmful BSD biosecurity multi-choice questions. We use the standard MCQ prompt format from the Inspect library (AI Security Institute, 2024).
Baselines. To obtain strong misuse uplift baselines, we evaluate several attacks [^3] We make the attacks, designed for open-ended tasks, compatible with multi-choice questions by jailbreaking the question and appending the answer choices to the attacking prompt.. First, we consider three single-turn jailbreaks—simple adaptive attacks (Andriushchenko et al., 2025), PAIR (Chao et al., 2025), and adversarial reasoning attacks (Sabbaghi et al., 2025)—which attempt to obtain an answer via a single model interaction. We also use Crescendo (Russinovich et al., 2024), a multi-turn jailbreak, and decomposition attacks (Jones et al., 2024, Glukhov et al., 2024), as described in Section˜3.2. Moreover, we include a new decomposition attack variant, described in detail in Section˜6. Each attack is run for five epochs (when applicable) with a budget of twelve calls to the strong model per task. Each row denotes a different strong model, and we use Qwen2.5-7B as (a) the attacker for PAIR and adversarial reasoning attacks and (b) the weak model for decomposition attacks. We also provide decomposition attack results with additional stronger open-weights models in Table˜3. We further provide baseline details in Appendix˜F.
Results. Across target models, decomposition attacks are highly effective and are often stronger than state-of-the-art jailbreaks. The effectiveness of decomposition attacks is most pronounced in the safest models. In Table˜1 and Table˜3, Claude‑3.5‑Sonnet and GPT‑5 perform poorly (have low misuse rates) under direct querying, consistent with the fact that they refuse most BSD questions in the direct setting. For these models, decomposition is particularly effective because it replaces a single obviously harmful request with a sequence of benign-looking sub-queries that the model is more likely to answer. Consistent with this mechanism, decomposition attack success correlates strongly with sub-query compliance: Claude‑3.5 answers 42.0% of sub-tasks despite answering only 3% of the original questions directly, and GPT‑5 answers 48.3% of sub-tasks despite only 1% direct answers. For target models with weaker safety filtering, jailbreak prompts are less likely to be blocked and thus achieve higher misuse rates. We study the detectability of decomposition versus jailbreak prompts in Section˜5.
Table 1: Misuse rate on BSD of attacks on various strong models. The performance of our decomposition pipeline (denoted by "New", see Section˜6 on our finetuning pipeline) on misuse uplift significantly increases when the decomposer is fine-tuned to produce better sub-queries, despite still lacking the requisite knowledge to solve the difficult BSD tasks.
@@
Direct query Adaptive PAIR Adversarial Reasoning Crescendo Decomposition Attack Decomposition Attack (New)
Claude‑3.5‑Sonnet 3.0 $\pm$ 0.2 27.3 $\pm$ 2.7 35.3 $\pm$ 2.4 46.7 $\pm$ 2.5 20.7 $\pm$ 2.1 41.6 $\pm$ 2.1 46.0 $\pm$ 2.4
Claude‑3.7‑Sonnet 15.0 $\pm$ 2.0 67.3 $\pm$ 3.0 62.7 $\pm$ 2.8 65.3 $\pm$ 2.5 52.7 $\pm$ 2.9 52.8 $\pm$ 2.2 65.6 $\pm$ 2.0
GPT‑4o 42.0 $\pm$ 3.2 42.0 $\pm$ 3.2 64.7 $\pm$ 2.7 70.7 $\pm$ 2.5 65.3 $\pm$ 2.8 68.4 $\pm$ 2.2 74.0 $\pm$ 2.0
o3‑mini 77.7 $\pm$ 1.9 84.7 $\pm$ 2.3 84.7 $\pm$ 2.0 84.0 $\pm$ 2.0 86.1 $\pm$ 2.2 82.0 $\pm$ 2.0 81.2 $\pm$ 2.3
o3 31.3 $\pm$ 2.0 32.7 $\pm$ 2.8 46.0 $\pm$ 2.4 56.7 $\pm$ 3.0 53.3 $\pm$ 2.6 52.0 $\pm$ 2.1 68.8 $\pm$ 2.0
Gemini-2.5-pro 64.7 $\pm$ 2.0 88.7 $\pm$ 1.7 88.0 $\pm$ 1.6 88.7 $\pm$ 2.0 86.0 $\pm$ 1.6 79.3 $\pm$ 2.2 82.0 $\pm$ 2.1
GPT-5 1.3 $\pm$ 1.0 1.3 $\pm$ 1.0 13.3 $\pm$ 1.6 18.7 $\pm$ 2.0 13.0 $\pm$ 1.2 45.3 $\pm$ 2.5 50.6 $\pm$ 2.2
@@ 5 Detectability and defense
Figure 4: Decompositions are harder to identify than jailbreaks per-input. (Left) Hidden-layer UMAP projection of Llama-3-8B. Jailbreaking prompts (from Table˜1) fall near the refused BSD questions, whereas decompositions are mostly located near benign WMDP questions. A larger $n$, i.e. more decompositions, leads to queries that look more benign. (Middle/Right) Decompositions are significantly harder to classify than jailbreaks (see 5.1).
Real-world LLM misuse (Nimmo et al., 2025, Lebedev et al., 2025) typically involves completing multiple tasks, many of which may appear benign in isolation. This threat model is in tension with standard defenses, which assume that a harmful request is confined to a single context window. Decomposition attacks (Glukhov et al., 2024, Jones et al., 2024) exploit this oversight and systematically avoid detection by splitting a harmful task into benign subtasks. We find that decomposition attacks are much harder to detect than standard jailbreaks. In Figure˜4 (left), we plot Llama-3-8B activations on refused prompts (questions from the BSD dataset) and answered questions (WMDP), along with jailbreaks and decomposition subtasks. This provides evidence that jailbreaks fall closer to the refused BSD questions, while decompositions designed to appear benign are grouped with the answered WMDP questions. In the following, we find that this holds more generally across defenses.
Figure 5: Detecting decompositions attacks is difficult, especially when adversaries dilute harmful queries with benign ones (‘fraction harmful,’ see Section˜5.2). Rows: Top row shows PR curves for detecting jailbreaks (PAIR), bottom row shows decomposition attacks. Columns: Left column shows per-sequence classification, right column shows the buffer defense. Our user-level buffer defense (right column), which maintains context across user queries, consistently outperforms per-sequence classification (left column).
@@ 5.1 Detecting harmful inputs
We frame detection as deciding whether a given decomposition or jailbreak prompt was generated based on a query from WMDP or from BSD. Because frontier models almost never refuse WMDP (Li et al., 2024a) questions (see Section˜4), but consistently refuse BSD questions, we label WMDP-derived prompts as safe and BSD-derived prompts as unsafe. unsafe queries are decompositions of the refused BSD questions, while safe queries are WMDP decompositions that frontier models answer without refusal. We evaluate detectability as a binary classification problem and report ROC curves and their AUC that show how well a detector ranks unsafe over safe prompts. Because the performance depends on the rate of attacks (what fraction is harmful), we also show precision–recall (PR) curves and vary the malicious:benign ratio.
@@
As a baseline, we use Llama-Guard-3-8B to classify each prompt as safe or unsafe. In Figure˜4 (middle), we observe that for decomposition prompts, performance is only marginally better than random (AUC =0.60=0.60). In contrast, the same detector is significantly more effective on jailbreak prompts (AUC =0.71=0.71).
Hardening LlamaGuard against decomposition attacks.
To evaluate the attacks under favourable conditions for the defender, we provide them with 20% of both the safe (from the WMDP dataset) and unsafe (from our BSD) decomposition or jailbreak prompts, where all attacks are generated by Qwen-2.5 7B. We finetune Llama-Guard-3-8B with the defender split, and calibrate the model to the new class distribution with Platt scaling. Adversarial training raises the AUC of Llama-Guard-3-8B for the jailbreak prompts (AUC =0.89=0.89) and the decompositions (AUC =0.76=0.76). So, while training with the exact attack distribution improves the defense, detection for decomposition attacks is difficult.
@@ 5.2 Detecting harmful users
Input-output detectors struggle to detect decomposition attacks because they lack the context necessary to recognize a harmful task distributed into multiple benign queries (see Section 5.1). We therefore lift the task to the user level: the defender succeeds if they flag an attacker before they are able to complete a harmful task. This assumes the model provider can trace all queries back to a single user, even if they occur in different sessions. We discuss the assumption in Section˜B.2. We will first show that an adaptive attack breaks a simple baseline defense in this setting; we will use this to motivate our buffer defense.
Rolling-window defense.
We first consider a simple baseline, where a defender inspects the last $k$ queries in a user’s history to identify potential misuse. $k$ is limited by the context length of a classifier, so $k<100$. In Figure˜5, this defense works well when the attacker queries with mostly harmful prompts (e.g. 1 in 2 prompts are harmful); however, it fails with the following adaptive attack.
An adaptive attack against buffer defenses.
Anticipating that the defender is using a buffer (or is tuning their classification threshold), the attacker mixes in a large number of ‘distractor’ benign queries. As the benign-to-malicious ratio increases (lighter traces in Figure˜5, left), classifier precision falls. In the case of a rolling window buffer, the detector’s fixed context is overflowed by innocuous messages, pushing earlier harmful steps out of context. See Section˜F.4 for more discussion.
@@
Instead of keeping a rolling-window, for every user we maintain the $m$ prompts (out of a total of $q$) with the highest prompt-wise harmful scores in a buffer (as scored prompt-wise by the adversarially trained Llama-Guard discussed above). For a new query, the buffer (ordered from most to least harmful) is fed in-context to either the adversarially trained Llama-Guard-3-8B or Llama-3-70B, which labels the user as either harmful or benign. The user is flagged if harmful is returned and subsequent requests are denied. Because the stateful buffer caches the $m$ most suspicious prompts across a user’s history, as opposed to the naive defense, it achieves non-trivial precision even when queries with a large number of distractor benign prompts.
To evaluate this setting, we randomly sample a benign or harmful user, and sample $q\in\{10,\ldots,400\}$ previous user queries; we use a buffer size of $m=\min(q,30)$. If the user is benign, then the correct label of all prompts is benign. If the user is unsafe, there is some probability (the benign-to-malicious ratio) that a prompt is harmful. We rebuild the buffer with per-prompt harmfulness scores at each draw, place the buffer in-context to the classifier, and prompt it to reason over the previous prompts in the buffer to identify patterns of misuse. We show precision-recall results for classifying over ‘users’ in Figure˜5 (right).
Results. Detection scores improve in the stateful setting. Stronger models with better in-context learning perform better in this setting. While user-level buffering only improves AUC from 0.76 to 0.79 for the fine-tuned Llama-Guard-3-8B, for Llama-3-70B the AUC improves from 0.71 to 0.86 This trend is clear in the right column of Figure˜5, where the precision–recall curve is shifted upward. Because our classifier requires only storing a few recent queries and uses a single additional inference call, it is relatively inexpensive. However, as can be seen in Figure˜5, the adaptive attack (˜An adaptive attack against buffer defenses.) can still drive the false-positive rate up significantly for all the classifiers we consider.
@@ 6 Scaling decomposition attacks
Figure 6: (Left, BSD) In the refusal setting, decomposition accuracy improves with more decompositions, while the baseline gets no answer from the strong model (GPT-4.1). (Right, WMDP) Decompositions underperform direct querying for benign tasks, suggesting the success of decomposition scaling is not from general test-time compute scaling.
We formally describe the decomposition attack in Section˜B.4. The success of a decomposition attack depends on the quality of generated sub-queries, which, in turn, depends on factors including the coarseness of the decomposition and how comprehensively they span the original task. We improve its performance over previous work (Jones et al., 2024) with two modifications: increasing the number of sub-tasks and distilling the model performing the decomposition.
Inference scaling. We study decomposition attack performance when we scale sub-tasks. In Figure˜6 (left), we sweep the number of decopmositions, using Mixtral-8x22B as the weak model and GPT-4.1 as the strong model. We find that accuracy improves as decompositions increase. In contrast, decomposition attacks are not generally useful for test-time scaling. Where the target is willing to answer directly, Figure˜6 (right), decomposition attacks harm performance. Attack gains are therefore due to uplift from the strong model’s answers.
Distilling an unsafe decomposer model with benign data. We improve decomposition quality on harmful misuse tasks by training the model to perform better decompositions through benign‑only tasks. Thus, we only fine‑tune the Decomposer model. We collect 700 teacher‑generated decompositions for (benign) MMLU-auxiliary (Hendrycks et al., 2020) questions using o3‑mini, then fine-tune Qwen on these prompts. For the distillation data, we randomly choose o3-mini to decompose a given question into 3, 6, or 12 parts. Restricting the process to benign prompts allows us to use stronger models for distillation regardless of their safety guardrails. Fine-tuning details are in Appendix˜F. Furthermore, ablations in Section˜F.3 show the performance gains come from improved decomposition quality, not from broader capability increases due to fine-tuning.
The results of the modified algorithm—deploying the fine-tuned model with increased number of decompositions—are in Table˜1. As the last two columns show, our method achieves the state-of-the-art on Claude-3.5, OpenAI-o3, GPT-4o, and GPT-5. Additionally, in Appendix˜G, we discuss that adding a jailbreaking method such as GCG Zou et al. (2023) helps to further increase the compliance of open-box models like Llama-3.1-8B and subsequently improve the misuse uplift.
@@ 7 Conclusion
We introduce a evaluation framework for measuring misuse uplift and detectability. Whereas previous evaluations measure if an attack can elicit harm from a given model, our framework measures the extent to which a strong model aides in misuse. We construct a threat model with realistic affordances for both the attacker (the ability to use weaker models) and the defender (tracking user queries across independent user conversations to detect misuse across contexts). We find that decomposition attacks (Jones et al., 2024, Glukhov et al., 2024) are a particularly effective attack in this setting, outperforming state-of-the-art single- and multi-turn jailbreaks. We develop a defense that mitigates misuse with a stateful monitor that reasons over many independent user inputs to detect clusters of harmful inputs; however, we find that decomposition attacks can still often subvert such detectors.
Ethics Statement
We build a dataset of biology and cybersecurity questions that violate the content restrictions of various frontier model providers. This is because realistic misuse uplift evaluations require questions that are both difficult and consistently refused. For example, we found that tasks that were not dual-use, e.g., difficult math questions (Rein et al., 2024) or standard jailbreaking behaviors (Mazeika et al., 2024) failed to yield any questions that met our difficult and refusal criteria (Section˜4). Given concerns around disseminating harmful information, we do not do a full release of the BSD questions. Instead, we commit to a restricted release of BSD under controlled access only.
This research studies and develops strong attacks to misuse language models. While presenting these methods could enable attackers, suppressing them would likely hurt progress on effective countermeasures. Consistent with security-through-transparency norms, we discuss both attacks and mitigation strategies (Sections 45). We maintain that the security benefits of empowering the research community outweigh the incremental risk of adversary adoption.